# Multinucleon transfer with time-dependent covariant density functional theory

### 12 Jan 2024 v1

@@ I Introduction
Numerous theoretical models have been developed and applied in studies of multinucleon transfer reactions, including the GRAZING model Winther (1994, 1995), the dinuclear system (DNS) model Niu et al. (2017); Feng (2017); Zhu et al. (2018); Guo et al. (2019); Bao (2020); Zhang et al. (2023); Liao et al. (2023), Langevin equations Zagrebaev and Greiner (2008a, b, 2011, 2013); Karpov and Saiko (2017); Saiko and Karpov (2019), the improved quantum molecular dynamics (ImQMD) model Zhao et al. (2015); Li et al. (2016); Zhu et al. (2017); Li et al. (2019); Zhao et al. (2021, 2022); Liao et al. (2023), and the time-dependent Hartree-Fock (TDHF) theory Simenel (2010); Sekizawa and Yabana (2013, 2014, 2016); Sekizawa (2017a, b); Wu and Guo (2019); Jiang and Wang (2020). Nuclear density functional theory (DFT), in particular, successfully reproduces and predicts basic properties for most nuclides on the nuclear chart Vautherin and Brink (1972); Dechargé and Gogny (1980); Ring (1996); Vretenar et al. (2005); Nikšić et al. (2011); Meng (2016). Its dynamical extension, the time-dependent density functional theory (TD-DFT), presents a general microscopic framework that, without any parameter specifically adjusted to the reaction mechanism, can be employed to explore multinucleon transfer reactions. Using a particular implementation of this framework, the TDHF, significant progress has been made in the study of multinucleon transfer reaction dynamics Simenel (2010); Sekizawa and Yabana (2013, 2014, 2016); Sekizawa (2017a, b); Wu and Guo (2019); Jiang and Wang (2020). Various factors related to the entrance channel properties such as neutron-to-proton ($N/Z$) ratio Sekizawa and Yabana (2013), the relative orientation of deformed ions Sekizawa and Yabana (2013, 2016); Sekizawa (2017a); Wu and Guo (2019), and the charge product $Z_{P}Z_{T}$ Sekizawa and Yabana (2013), have been found to affect the outcome of these reactions. Additionally, the inclusion of beyond-mean-field effects in methods such as in the time-dependent random-phase approximation (TDRPA) Williams et al. (2018); Godbey et al. (2020), and the stochastic mean-field theory (SMF) Ayik (2008); Ayik et al. (2020); Sekizawa and Ayik (2020); Ayik et al. (2021, 2023a, 2023b); Arik et al. (2023), has greatly enhanced the accuracy with which various observables in MNT reactions can be computed.
@@ III Numerical Details
In this work, we analyze the following reactions: ${}^{40}{\rm Ca}+{}^{124}{\rm Sn}$ at $E_{\rm lab}=170$ MeV, ${}^{40}{\rm Ca}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=249$ MeV, ${}^{58}{\rm Ni}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=328.4$ MeV, ${}^{40}{\rm Ar}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=256$ MeV, and ${}^{206}{\rm Pb}+{}^{118}{\rm Sn}$ at $E_{\rm lab}=1200$ MeV. The relativistic density functionals PC-PK1 Zhao et al. (2010) and DD-ME2 Lalazissis et al. (2005) are employed in the TD-CDFT calculation. Before the collision, the projectile and target are in their ground states obtained from stationary self-consistent mean-field calculations, with a grid size of $N_{x}\times N_{y}\times N_{z}=26\times 26\times 26$. The mesh spacing along each axis is $0.8$ fm. In the dynamical case, the grid size is extended to $N_{x}\times N_{y}\times N_{z}=60\times 26\times 60$. The collision takes place in the $x$-$z$ plane. For the reaction ${}^{206}{\rm Pb}+{}^{118}{\rm Sn}$, we have extended the space volume by setting the mesh spacing along each axis to $1.0$ fm. For the time evolution of single-particle wave functions, a predictor-corrector method is employed, with a fourth-order Taylor expansion of the time-evolution operator. The time step is $0.2$ fm/$c$.
@@ IV Results and discussion
Figure 1: (Color online) Cross sections for the transfer channels to the light fragment in the reaction ${}^{40}$a + ${}^{124}$n at $E_{\text{lab}}$ = 170 MeV. Each panel shows the cross section for a different proton-transfer channel, as a function of the PLF average neutron number. Black filled circles denote experimental results from Ref. Corradi et al. (1996). Red (blue) histograms represent results obtained by the TD-CDFT calculation with the density functional PC-PK1 (DD-ME2), while grey (shaded) histograms correspond to TDHF results from Ref. Sekizawa and Yabana (2013).
@@
Figure 1 displays the cross sections for proton transfer channels to the light fragment in the ${}^{40}$a + ${}^{124}$n reaction at $E_{\text{lab}}$ = 170 MeV, from one proton added $(+1p)$ to six protons removed $(-6p)$, as functions of the average number of neutrons in the PLF. The terms pick-up and stripping are conventionally referred to the lighter partner of the reaction. The data are from Ref. Corradi et al. (1996), and the red (blue) histograms represent results obtained by the TD-CDFT calculation with the density functional PC-PK1 (DD-ME2). For comparison, we also include the cross sections calculated using the TDHF model of Ref. Sekizawa and Yabana (2013), with the Skyrme functional SLy5 Chabanat et al. (1998). The largest value of the impact parameter for which fusion takes place is $3.93$ fm in the PC-PK1 calculation, which is close to value $3.95$ fm reported in Ref. Sekizawa and Yabana (2013). For the calculation with DD-ME2 this value is $4.24$ fm.
@@
As shown in the figure, the data for the $(0p)$ channel and $(-1p)$ channel are reproduced quite well by all three model calculations. For the $(-2p)$ channel, zero- to four-neutron pick-up channels, in which ${}^{40}$a captures between zero and four neutrons from ${}^{124}$n, are reproduced by theory. The reverse process, however, is underestimated by all three models. Before collision, the neutron-to-proton ratios $N/Z$ for the projectile and target nuclei are $1$ and $1.48$, respectively. One expects that, for collisions between nuclei with different $N/Z$ values, the dominant transfer process is towards charge equilibrium, that is, nucleon transfer tends to equalize the $N/Z$ ratio in the PLF and TLF. Here, this means neutron transfer from ${}^{124}$n to ${}^{40}$a, and proton transfer from Ca to Sn. Nucleons are transferred in the interval of impact parameters close the fusion critical impact parameter. As shown in more detail in Ref. Sekizawa and Yabana (2013), the average number of transferred nucleons decreases rapidly as the impact parameter increases. From $(-3p)$ to $(-6p)$, the calculated cross sections decrease faster than the experimental values, especially the ones obtained with the two relativistic density functionals. In addition, the peak positions of the theoretical cross sections are shifted to larger average neutron numbers with respect to data. As pointed out in previous studies Sekizawa and Yabana (2013), this is most probably due to the fact that the theoretical methods used to calculate cross sections do not properly take into account neutron evaporation effects. The process of neutron evaporation will naturally shift the cross sections to smaller values of the average neutron number of the PLF, that is, in better agreement with the observed values. We also notice that the cross sections calculated with the three density functionals start to differ significantly only for channels in which a larger number of protons is removed from the projectile.
@@
Similar results are obtained for the reaction ${}^{40}$a + ${}^{208}$b at $E_{\text{lab}}$ = 249 MeV. The theoretical cross sections obtained in the present calculation, and in Ref. Sekizawa and Yabana (2013) with the Skyrme functional SLy5, are compared with available data Szilner et al. (2005) in Fig. 2. The largest value of the impact parameter for which fusion occurs is $4.63$ fm for the functional PC-PK1, and $4.89$ fm in the calculation with DD-ME2, while the value $4.55$ fm was obtained for Skyrme functional SLy5 in Ref. Sekizawa and Yabana (2013). The three functionals predict comparable cross sections for the proton stripping channels. In the $(0p)$ channel, the cross sections for the pick-up of one to four neutrons are well reproduced, while the data for the addition of five and six neutrons in the PLF are slightly overestimated. Generally, we find that the theoretical results tend to slightly overestimate neutron pick-up and underestimate the neutron stripping cross sections in the $(-1p)$, $(-2p)$, and $(-3p)$ channels. As explained above, this discrepancy can be attributed to the fact that nucleon transfer generally follows the direction of charge equilibrium. In the present case, the neutron-to-proton ratio of the projectile is $1$, while that of the target is $1.54$, slightly higher than the former value $1.48$ for ${}^{124}$n. The calculated cross sections for the $(-4p)$, $(-5p)$, and $(-6p)$ channels are considerably reduced, compared to the data. This discrepancy arises from the fusion process occurring before the multi-proton stripping channels are included.
@@
Figure 4: (Color online) Cross sections for the transfer channels to the lighter fragment in the reaction ${}^{40}$r + ${}^{208}$b reaction at $E_{\text{lab}}$ = 256 MeV. Each panel shows the cross section for a different proton- transfer channel, as a function of the PLF average neutron number. Black filled circles denote experimental results from Ref. Mijatović et al. (2016). Red (blue) histograms denote results obtained in the TD-CDFT calculation with the density functional PC-PK1 (DD-ME2), while the grey (shaded) histograms correspond to GRAZING model calculation Winther (1994, 1995).
Figure 4 displays the cross sections from the $(+4p)$ to the $(-4p)$ channel, for the reaction ${}^{40}$r + ${}^{208}$b at $E_{\text{lab}}$ = 256 MeV. In this reaction, the neutron-to-proton ratio of the projectile is $1.22$, much closer to that of the target ($1.54$) than in the three previous examples. For the functional PC-PK1, the maximum impact parameter for which fusion occurs is $5.99$ fm, while it is $6.36$ fm for DD-ME2.
In the $(0p)$ channel, the calculation is in excellent agreement with the data, both in the neutron pick-up and neutron stripping channels. For the $(-1p)$ channel, the theoretical results slightly underestimate the experimental cross sections on the neutron stripping side, and overestimate the data for neutron pick-up. As seen in the $(-2p)$, $(-3p)$, and $(-4p)$ channels, this discrepancy gradually becomes larger with increasing the number of protons removed from the projectile, because of neutron evaporation that is not taken into account in model calculations. The theoretical results for the $(+1p)$, $(+2p)$, $(+3p)$, and $(+4p)$ channels generally show a very good agreement with the experimental values. In the case of neutron pick-up, however, there is a tendency to overestimate the neutron transfer in the $(+3p)$ and $(+4p)$ channels, which can be attributed to quasi-fission processes involving the transfer of a large number of nucleons. In our calculation, a rapid increase in the number of transferred nucleons is observed as the impact parameter decreases. Hence, for impact parameters smaller than 7 fm, we reduced the interval of $b$ values from $0.25$ fm to $0.05$ fm. Notably, when using the DD-ME2 functional, the $(+3p)$ and $(+4p)$ channels exhibit a distinct peak associated with quasi-fission, indicating a sudden change in reaction dynamics. When compared to results of the GRAZING model calculation Winther (1994, 1995), the data for the proton pick-up channels are much better reproduced by the TD-CDFT, especially for the $(+2p)$, $(+3p)$ and $(+4p)$ channels.
@@
The calculated cross sections for the transfer channels from $(+1p)$ to $(-2p)$ to the lighter fragment in the reaction ${}^{206}$b + ${}^{118}$n at $E_{\text{lab}}$ = 1200 MeV, are compared with the data from Ref. Diklić et al. (2023), and to GRAZING model calculation in Fig. 5. The incident energy is below the fusion barrier. In the $(-2p)$ channel, the magnitude of the cross section is comparable to the data. The position of the peak is at $N=67$, which is three neutrons more than in the experimental data. The neutron-to-proton ratio of the projectile is $1.51$, larger than that of the target at $1.36$. The direction of charge equilibrium, therefore, is the transfer of neutrons from the projectile to the target, while protons are transferred in the opposite direction. The position of the peak in the $(-1p)$ and $(0p)$ channels coincides with the experimental data. For the $(+1p)$ channel, the PC-PK1 and DD-ME2 results are in better agreement with the data compared to the GRAZING model. Even though the peak positions are predicted at larger neutron number, overall the calculated cross sections reproduce the experimental data to a very good level of agreement.
@@ V Summary
The time-dependent CDFT with particle number projection has been applied to a study of multinucleon transfer reactions. By employing two relativistic density functionals PC-PK1 and DD-ME2, we have calculated the MNT total cross sections for five reactions. The results for ${}^{40}{\rm Ca}+{}^{124}{\rm Sn}$ at $E_{\rm lab}=170$ MeV, ${}^{40}{\rm Ca}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=249$ MeV, and ${}^{58}{\rm Ni}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=328.4$ MeV, have been compared to available data, and theoretical cross sections obtained in a previous study based on the Skyrme functional SLy5. In a sense these are benchmark examples, here used to verify our model and the particular numerical implementation.
For all three reactions, in the $(0p)$ and $(-1p)$ channels, the calculated cross sections are found in very good agreement with the data. In the $(-2p)$ and $(-3p)$ channels, the results generally reproduce the magnitude of the experimental cross sections, but tend to underestimate neutron stripping channels. In the $(-4p)$, $(-5p)$ and $(-6p)$ channels, the theoretical cross sections decrease much more rapidly than the data, and the calculated peak positions are shifted to larger average neutron numbers with respect to the experimental peaks. It is generally expected that, when neutron evaporation effects are taken into account, the peak positions will shift towards smaller average neutron numbers, providing a better description of the data. An interesting result is that the three density functionals: Skyrme SLy5, the relativistic point-coupling PC-PK1, and the finite-range meson exchange DD-ME2, predict similar MNT cross sections for all three reactions, even though the effective interactions are very different and their parameters were adjusted to nuclear ground-state properties following vastly different protocols. This demonstrates the robustness of the microscopic approach, based on nuclear density functionals, in the description of reaction dynamics.

### 22 Feb 2024 v2

@@ I Introduction
Numerous theoretical models have been developed and applied in studies of multinucleon transfer reactions, including the GRAZING model Winther (1994, 1995); Wen et al. (2019), the dinuclear system (DNS) model Feng (2017); Niu et al. (2017); Zhu et al. (2017); Bao et al. (2018); Zhu et al. (2018); Guo et al. (2019); Zhang et al. (2023); Liao et al. (2023a, b); Zhao et al. (2023); Feng (2023), Langevin equations Zagrebaev and Greiner (2008a, b, 2011, 2013); Karpov and Saiko (2017); Saiko and Karpov (2019), the improved quantum molecular dynamics (ImQMD) model Zhao et al. (2015); Li et al. (2016); Yao and Wang (2017); Li et al. (2019); Zhao et al. (2021, 2022); Liao et al. (2023b), and the time-dependent Hartree-Fock (TDHF) theory Simenel (2010); Sekizawa and Yabana (2013, 2014, 2016); Sekizawa (2017a, b); Wu and Guo (2019); Jiang and Wang (2020); Wu et al. (2022). Nuclear density functional theory (DFT), in particular, successfully reproduces and predicts basic properties for most nuclides on the nuclear chart Vautherin and Brink (1972); Dechargé and Gogny (1980); Ring (1996); Vretenar et al. (2005); Nikšić et al. (2011); Meng (2016). Its dynamical extension, the time-dependent density functional theory (TD-DFT), presents a general microscopic framework that, without any parameter specifically adjusted to the reaction mechanism, can be employed to explore multinucleon transfer reactions. Using a particular implementation of this framework, the TDHF, significant progress has been made in the study of multinucleon transfer reaction dynamics Simenel (2010); Sekizawa and Yabana (2013, 2014, 2016); Sekizawa (2017a, b); Wu and Guo (2019); Jiang and Wang (2020); Wu et al. (2022). Various factors related to the entrance channel properties such as neutron-to-proton ($N/Z$) ratio Sekizawa and Yabana (2013), the relative orientation of deformed ions Sekizawa and Yabana (2013, 2016); Sekizawa (2017a); Wu and Guo (2019), and the charge product $Z_{P}Z_{T}$ Sekizawa and Yabana (2013), have been found to affect the outcome of these reactions. Additionally, the inclusion of beyond-mean-field effects in methods such as in the time-dependent random-phase approximation (TDRPA) Williams et al. (2018); Godbey et al. (2020), and the stochastic mean-field theory (SMF) Ayik (2008); Ayik et al. (2020); Sekizawa and Ayik (2020); Ayik et al. (2021, 2023a, 2023b); Arik et al. (2023), has greatly enhanced the accuracy with which various observables in MNT reactions can be computed.
@@ III Numerical Details
In this work, we analyze the following reactions: ${}^{40}{\rm Ca}+{}^{124}{\rm Sn}$ at $E_{\rm lab}=170$ MeV, ${}^{40}{\rm Ca}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=249$ MeV, ${}^{58}{\rm Ni}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=328.4$ MeV, ${}^{40}{\rm Ar}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=256$ MeV, and ${}^{206}{\rm Pb}+{}^{118}{\rm Sn}$ at $E_{\rm lab}=1200$ MeV. The relativistic density functionals PC-PK1 Zhao et al. (2010) and DD-ME2 Lalazissis et al. (2005) are employed in the TD-CDFT calculation. Before the collision, the projectile and target are in their ground states obtained from stationary self-consistent mean-field calculations Ren et al. (2017); Zhang et al. (2022); Zhang (2023), with a grid size of $N_{x}\times N_{y}\times N_{z}=26\times 26\times 26$. The mesh spacing along each axis is $0.8$ fm. In the dynamical case, the grid size is extended to $N_{x}\times N_{y}\times N_{z}=60\times 26\times 60$. The collision takes place in the $x$-$z$ plane. For the reaction ${}^{206}{\rm Pb}+{}^{118}{\rm Sn}$, we have extended the space volume by setting the mesh spacing along each axis to $1.0$ fm. For the time evolution of single-particle wave functions, a predictor-corrector method is employed, with a fourth-order Taylor expansion of the time-evolution operator. The time step is $0.2$ fm/$c$.
@@ IV Results and discussion
Figure 1: (Color online) Cross sections for the transfer channels to the light fragment in the reaction ${}^{40}$a + ${}^{124}$n at $E_{\text{lab}}$ = 170 MeV. Each panel shows the cross section for a different proton-transfer channel, as a function of the PLF neutron number. Black filled circles denote experimental results from Ref. Corradi et al. (1996). Red (blue) histograms represent results obtained by the TD-CDFT calculation with the density functional PC-PK1 (DD-ME2), while grey (shaded) histograms correspond to TDHF results from Ref. Sekizawa and Yabana (2013).
@@
Figure 1 displays the cross sections for proton transfer channels to the light fragment in the ${}^{40}$a + ${}^{124}$n reaction at $E_{\text{lab}}$ = 170 MeV, from one proton added $(+1p)$ to six protons removed $(-6p)$, as functions of the number of neutrons in the PLF. The terms pick-up and stripping are conventionally referred to the lighter partner of the reaction. The data are from Ref. Corradi et al. (1996), and the red (blue) histograms represent results obtained by the TD-CDFT calculation with the density functional PC-PK1 (DD-ME2). For comparison, we also include the cross sections calculated using the TDHF model of Ref. Sekizawa and Yabana (2013), with the Skyrme functional SLy5 Chabanat et al. (1998). The largest value of the impact parameter for which fusion takes place is $3.93$ fm in the PC-PK1 calculation, which is close to value $3.95$ fm reported in Ref. Sekizawa and Yabana (2013). For the calculation with DD-ME2 this value is $4.24$ fm.
@@
As shown in the figure, the data for the $(0p)$ channel and $(-1p)$ channel are reproduced quite well by all three model calculations. For the $(-2p)$ channel, zero- to four-neutron pick-up channels, in which ${}^{40}$a captures between zero and four neutrons from ${}^{124}$n, are reproduced by theory. The reverse process, however, is underestimated by all three models. Before collision, the neutron-to-proton ratios $N/Z$ for the projectile and target nuclei are $1$ and $1.48$, respectively. One expects that, for collisions between nuclei with different $N/Z$ values, the dominant transfer process is towards charge equilibrium, that is, nucleon transfer tends to equalize the $N/Z$ ratio in the PLF and TLF. Here, this means neutron transfer from ${}^{124}$n to ${}^{40}$a, and proton transfer from Ca to Sn. Nucleons are transferred in the interval of impact parameters close the fusion critical impact parameter. The mean values of the primary fragment distribution $P_{N,Z}$, for each impact parameter approach the equilibrium charge asymmetry values. At low energies, the mean-field approximation provides a good description for the most probable path, with very small dispersion of charge and mass distributions of primary fragments. In addition, the product form of $P_{N,Z}$ in Eq.(9) indicates that there are no correlations between neutron and proton distributions. As shown in more detail in Ref. Sekizawa and Yabana (2013), the average number of transferred nucleons decreases rapidly as the impact parameter increases. From $(-3p)$ to $(-6p)$, the calculated cross sections decrease faster than the experimental values, especially the ones obtained with the two relativistic density functionals. In addition, the peak positions of the theoretical cross sections are shifted to larger average neutron numbers with respect to data. As pointed out in previous studies Sekizawa and Yabana (2013), this is most probably due to the fact that the theoretical methods used to calculate cross sections do not take into account neutron evaporation effects. The process of neutron evaporation will naturally shift the cross sections to smaller values of the average neutron number of the PLF, that is, in better agreement with the observed values. We also notice that the cross sections calculated with the three density functionals start to differ significantly only for channels in which a larger number of protons is removed from the projectile.
@@
Similar results are obtained for the reaction ${}^{40}$a + ${}^{208}$b at $E_{\text{lab}}$ = 249 MeV. The theoretical cross sections obtained in the present calculation, and in Ref. Sekizawa and Yabana (2013) with the Skyrme functional SLy5, are compared with available data Szilner et al. (2005) in Fig. 2. The largest value of the impact parameter for which fusion occurs is $4.63$ fm for the functional PC-PK1, and $4.89$ fm in the calculation with DD-ME2, while the value $4.55$ fm was obtained for Skyrme functional SLy5 in Ref. Sekizawa and Yabana (2013). The three functionals predict comparable cross sections for the proton stripping channels. In the $(0p)$ channel, the cross sections for the pick-up of one to four neutrons are well reproduced, while the data for the addition of five and six neutrons in the PLF are slightly overestimated. Generally, we find that the theoretical results tend to slightly overestimate neutron pick-up and underestimate the neutron stripping cross sections in the $(-1p)$, $(-2p)$, and $(-3p)$ channels. As explained above, this discrepancy can be attributed to the fact that nucleon transfer generally follows the direction of charge equilibrium. In the present case, the neutron-to-proton ratio of the projectile is $1$, while that of the target is $1.54$, slightly higher than the value $1.48$ for ${}^{124}$n in the previous example. The calculated cross sections for the $(-4p)$, $(-5p)$, and $(-6p)$ channels are considerably reduced, compared to the data. This discrepancy arises from the fusion process occurring before the multi-proton stripping channels are included.
@@
Figure 4: (Color online) Cross sections for the transfer channels to the lighter fragment in the reaction ${}^{40}$r + ${}^{208}$b reaction at $E_{\text{lab}}$ = 256 MeV. Each panel shows the cross section for a different proton- transfer channel, as a function of the PLF neutron number. Black filled circles denote experimental results from Ref. Mijatović et al. (2016). Red (blue) histograms denote results obtained in the TD-CDFT calculation with the density functional PC-PK1 (DD-ME2), while the grey (shaded) histograms correspond to GRAZING model calculation Winther (1994, 1995).
Figure 4 displays the cross sections from the $(+4p)$ to the $(-4p)$ channel, for the reaction ${}^{40}$r + ${}^{208}$b at $E_{\text{lab}}$ = 256 MeV. In this reaction, the neutron-to-proton ratio of the projectile is $1.22$, much closer to that of the target ($1.54$) than in the previous three examples. For the functional PC-PK1, the maximum impact parameter for which fusion occurs is $5.99$ fm, while this value is $6.36$ fm for DD-ME2.
In the $(0p)$ channel, the calculation is in excellent agreement with the data, both in the neutron pick-up and neutron stripping channels. For the $(-1p)$ channel, the theoretical results slightly underestimate the experimental cross sections on the neutron stripping side, and overestimate the data for neutron pick-up. As seen in the $(-2p)$, $(-3p)$, and $(-4p)$ channels, this discrepancy gradually becomes larger with increasing the number of protons removed from the projectile, because of neutron evaporation that is not taken into account in model calculations. The theoretical results for the $(+1p)$, $(+2p)$, $(+3p)$, and $(+4p)$ channels generally show a very good agreement with the experimental values. In the case of neutron pick-up, however, there is a tendency to overestimate the neutron transfer in the $(+3p)$ and $(+4p)$ channels, which can be attributed to quasi-fission processes involving the transfer of a large number of nucleons. In our calculation, a rapid increase in the number of transferred nucleons is observed as the impact parameter decreases. Hence, for impact parameters smaller than 7 fm, we reduced the interval of $b$ values from $0.25$ fm to $0.05$ fm. Notably, in the calculation with the functional DD-ME2, the $(+3p)$ and $(+4p)$ channels exhibit a distinct peak associated with quasi-fission, indicating a sudden change in reaction dynamics. When compared to results of the GRAZING model calculation Winther (1994, 1995), the data for the proton pick-up channels are much better reproduced by the TD-CDFT, especially for the $(+2p)$, $(+3p)$ and $(+4p)$ channels.
@@
The calculated cross sections for the transfer channels from $(+1p)$ to $(-2p)$ to the lighter fragment in the reaction ${}^{206}$b + ${}^{118}$n at $E_{\text{lab}}$ = 1200 MeV, are compared with the data from Ref. Diklić et al. (2023), and to GRAZING model calculation in Fig. 5. In the $(-2p)$ channel, the magnitude of the cross section is comparable to the data. The position of the peak is at $N=67$, which is three neutrons more than in the experimental data. The neutron-to-proton ratio of the projectile is $1.51$, larger than that of the target at $1.36$. The direction of charge equilibrium, therefore, is the transfer of neutrons from the projectile to the target, while protons are transferred in the opposite direction. The position of the peak in the $(-1p)$ and $(0p)$ channels coincides with the experimental data. For the $(+1p)$ channel, the PC-PK1 and DD-ME2 results are in better agreement with the data compared to the GRAZING model. Even though the peak positions are predicted at larger neutron number, overall the calculated cross sections reproduce the experimental data to a very good level of agreement.
@@ V Summary
The time-dependent CDFT with particle number projection has been applied to a study of multinucleon transfer reactions. By employing two relativistic density functionals PC-PK1 and DD-ME2, we have calculated the MNT total cross sections for five reactions. The results for ${}^{40}{\rm Ca}+{}^{124}{\rm Sn}$ at $E_{\rm lab}=170$ MeV, ${}^{40}{\rm Ca}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=249$ MeV, and ${}^{58}{\rm Ni}+{}^{208}{\rm Pb}$ at $E_{\rm lab}=328.4$ MeV, have been compared to available data, and theoretical cross sections obtained in a previous study based on the Skyrme functional SLy5. These are benchmark examples, here used to validate our model and the particular numerical implementation.
For all three reactions, in the $(0p)$ and $(-1p)$ channels, the calculated cross sections are found in very good agreement with the data. In the $(-2p)$ and $(-3p)$ channels, the results generally reproduce the magnitude of the experimental cross sections, but tend to underestimate neutron stripping channels. In the $(-4p)$, $(-5p)$ and $(-6p)$ channels, the theoretical cross sections decrease much more rapidly than the data, and the calculated peak positions are shifted to larger neutron numbers with respect to the experimental peaks. It is generally expected that, when neutron evaporation effects are taken into account, the peak positions will shift towards smaller neutron numbers. In general, model calculations reproduce the data for production cross-sections of primary fragments near the equilibrium charge asymmetry values. As primary products are removed further away from the equilibrium charge asymmetry values, the rate of decrease of the theoretical cross sections is much faster compared to data. In these channels, the calculated cross sections can be one or two orders of magnitude smaller than their experimental counterparts. An interesting result is that the three density functionals: Skyrme SLy5, the relativistic point-coupling PC-PK1, and the finite-range meson exchange DD-ME2, predict similar MNT cross sections for all three reactions, even though the effective interactions are very different and their parameters were adjusted to nuclear ground-state properties following vastly different protocols. This demonstrates the robustness of the microscopic approach, based on nuclear density functionals, in the description of reaction dynamics.